# On the Statistical Complexity of Estimating VENDI Scores from Empirical Data

### v1

@@ Abstract
Reference-free evaluation metrics for generative models have recently been studied in the machine learning community. As a reference-free metric, the VENDI score [1] quantifies the diversity of generative models using matrix-based entropy from information theory. The VENDI score is usually computed through the eigendecomposition of an $n\times n$ kernel matrix for $n$ generated samples. However, due to the high computational cost of eigendecomposition for large $n$, the score is often computed on sample sizes limited to a few tens of thousands. In this paper, we explore the statistical convergence of the VENDI score and demonstrate that for kernel functions with an infinite feature map dimension, the evaluated score for a limited sample size may not converge to the matrix-based entropy statistic. We introduce an alternative statistic called the $t$-truncated VENDI statistic. We show that the existing Nyström method and the FKEA approximation method for the VENDI score will both converge to the defined truncated VENDI statistic given a moderate sample size. We perform several numerical experiments to illustrate the concentration of the empirical VENDI score around the truncated VENDI statistic and discuss how this statistic correlates with the visual diversity of image data.
@@ 1 Introduction
The widespread use of generative artificial intelligence (AI) has underscored the need for accurate evaluation of generative models based on the quality and diversity of their generated samples. Users of generative AI often have access to multiple models trained using different algorithms and datasets, necessitating efficient evaluation mechanisms to choose the best model from the available options. The applicability of a model evaluation approach depends on factors such as the required generated sample size, computational cost, and the availability of reference data. Several recent studies on evaluating generative models have introduced assessment methods that relax the requirements for data and computational resources.
Specifically, to enable the evaluation of generative models in settings without reference data, the recent literature has focused on reference-free evaluation scores, which remain applicable in the absence of a reference dataset. The VENDI score [1] is one such reference-free metric that quantifies the diversity of generated data using the entropy of a kernel similarity matrix formulated for the generated samples. As analyzed by [1] and [2], the reference-free assessment of the VENDI score can be interpreted as an unsupervised identification of clusters within the generated data, followed by the calculation of the entropy of the detected cluster variable. Due to its flexibility and adaptability to various domains, the VENDI score has been applied to measure the diversity of samples across different modalities, including image, text, and video data.
@@
Figure 1: Statistical convergence of VENDI score for different sample sizes on ImageNet data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=40$. DINOv2 embedding (dimension 768) is used in computing the score.
While the VENDI score does not require reference samples, its computational cost increases rapidly with the number of generated samples $n$. In practice, calculating the entropy of the eigenvalues of the $n\times n$ kernel matrix involves performing an eigendecomposition, which requires $O(n^{3})$ computations. As a result, the computational load becomes substantial for a large sample size $n$, and the VENDI score is typically evaluated for sample sizes limited to a few tens of thousands. Consequently, the exact VENDI score, as defined in its original formulation, is usually not computed for sample sizes exceeding 20,000. A key question that arises is whether the VENDI score estimated from empirical data has converged to the underlying statistic, i.e., the limit value as the sample size approaches infinity. The statistical convergence has not been thoroughly investigated for models trained on large-scale datasets, e.g. ImageNet [3], which contain many sample categories and may require a large sample size for proper assessment.
@@
In this work, we study the statistical convergence of the VENDI score and aim to analyze the statistic estimated by the VENDI score for large-scale image, text, and video generative models. We discuss the answer to the convergence question for two types of kernel functions: 1) kernel functions with a finite feature dimension, e.g. the cosine similarity and polynomial kernels, 2) kernel functions with an infinite feature map such as Gaussian (RBF) and Laplace kernels. For kernel functions with a bounded feature dimension $d$, we theoretically and numerically show that a sample size $n=O(d)$ is sufficient to guarantee convergence to the VENDI statistic. For example, the left plot in Figure 1 shows that in the case of the cosine similarity kernel, the VENDI score on $n$ randomly selected ImageNet samples converges as the sample size reaches $5000$, where the feature dimension $d$ (using standard DINOv2 embedding) is 768.
@@
On the other hand, our numerical results for kernel functions with an infinite feature map suggest that in practical scenarios with diverse generative models and datasets, a sample size bounded by 20000 is insufficient for convergence to the VENDI statistic. For example, the right plot in Figure 1 shows the evolution of the VENDI score with Gaussian kernel on the ImageNet data, and the VENDI score keeps growing at a considerable rate even with 20,000 samples. As the costs of computing the exact score will be highly expensive beyond 20,000, it will be hard to empirically estimate the required sample size for convergence to the VENDI statistic.
Given the difference between the VENDI score for $n=O(10^{4})$ and the VENDI statistic (when $n\rightarrow\infty$) in the case of infinite kernel feature dimension, an important question is how to interpret the statistic estimated by the VENDI score when the sample size $n$ is bounded. We attempt to address the question by introducing an alternative statistic, which we call the $t$-truncated VENDI statistic. This statistic is calculated using only the top-$t$ eigenvalues of the kernel covariance matrix, excluding the remaining eigenvalues in the calculation of the VENDI statistic. We prove that a sample size $n=O(t)$ is enough to estimate the $t$-truncated VENDI statistic from $n$ generated samples whether the kernel feature dimension is finite or infinite. This result suggests the affordability of estimating the $t$-truncated VENDI statistic for every kernel function.
To connect the defined $t$-truncated VENDI statistic to existing computation methods for the VENDI score, we show that the standard methods proposed for approximating the VENDI score can be viewed as estimations of the $t$-truncated VENDI statistic. The approximation methods include the Nyström method as applied by [1] and the random Fourier feature-based FKEA method [4] for shift-invariant kernels. We show that both the Nyström and FKEA methods with dimension hyperparameter $t$ provide an estimation of the $t$-truncated VENDI statistic. Our results indicate that the truncated VENDI statistic is implicitly estimated by the computationally efficient methods proposed by [1] and [4].
@@
We perform several numerical experiments to validate the connection between the VENDI scores computed using a bounded $n$ samples and our defined $t$-truncated VENDI statistic. Our numerical results on several image, text, and video datasets and generative models indicate that in the case of a finite-dimension kernel map, the VENDI score efficiently converges to the VENDI statistic. On the other hand, in the case of infinite-dimension Gaussian kernel functions, we numerically observe the growth of the score beyond $n=$10,000. Our numerical results further confirm that the scores computed by Nyström method in [1] and the FKEA method [4] provide tight estimations of the truncated VENDI statistic which could provide a statically affordable substitute for the VENDI statistic in practical scenarios with large-scale datasets. The following is a summary of this work’s contributions:
  • Analyzing the statistical convergence of the reference-free VENDI score under finite and infinite kernel feature maps,
  • Providing numerical evidence on the gap between the VENDI score and VENDI statistic for infinite kernel feature maps with bounded sample size $n$,
  • Introducing the truncated VENDI statistic as a diversity-related statistic for generative models
  • Demonstrating convergence of the Nyström and FKEA methods’ proxy VENDI score to the truncated VENDI statistic.
@@ 2 Related Works
Diversity evaluation for generative models Diversity evaluation in generative models can be categorized into two primary types: reference-based and reference-free methods. Reference-based approaches rely on a predefined dataset to assess the diversity of generated data. Metrics such as FID [5] and KID [6] measure the distance between the generated data and the reference, while Recall [7, 8] and Coverage[9] evaluate the extent to which the generative model captures existing modes in the reference dataset. In contrast, the reference-free metrics, VENDI [1][10] and RKE [2], assign diversity scores based on the eigenvalues of a kernel similarity matrix of the generated data. [2]’s results interpret the approach as identifying modes and their frequencies within the generated data followed by entropy calculation for the frequency parameters. In this work, we specifically focus on the statistical convergence of the reference-free VENDI and RKE scores.
@@
Efficient computation of matrix-based entropy. Several strategies have been proposed in the literature to reduce the computational complexity of matrix-based entropy calculations, which involve the computation of matrix eigenvalues—a process that scales cubically with the size of the dataset. [11] propose an efficient algorithm for approximating matrix-based Renyi’s entropy of arbitrary order $\alpha$, which achieves a reduction in computational complexity down to $O(n^{2}sm)$ with $s,m\ll n$. Additionally, kernel matrices can be approximated using low-rank techniques such as incomplete Cholesky decomposition [12, 13] or CUR matrix decompositions [14], which provide substantial computational savings. [10] suggest to leverage Nyström method [15] with $m$ components, which results in $O(nm^{2})$ computational complexity. Further reduction in complexity is possible using Random Fourier Features, as suggested by [4], which allows the computation to scale linearly with $O(n)$ as a function of the dataset size. This work focuses on the latter two methods and which statistics they estimate.
@@
Impact of embedding spaces on diversity evaluation. In our image-related experiments, we used the DinoV2 embedding [16], as [17] demonstrate the alignment of this embedding with human evaluations. We note that the kernel function in the VENDI score can be similarly applied to other embeddings, including the standard InceptionV3[18] and CLIP embeddings [19] as suggested by [20]. Also, in our experiments on text data, we utilized the text-embedding-3-large [21] model, and for the video experiments, we employed the I3D embedding [22]. We use the mentioned embeddings in our experiments, while our theoretical results suggest that the convergence behavior would be similar for other embeddings.
@@ 3 Preliminaries
Consider a generative model $\mathcal{G}$ that generates samples from a probability distribution $P_{X}$. To conduct a reference-free evaluation of the model, we suppose the evaluator has access to $n$ independently generated samples from $P_{X}$, denoted by $x_{1},\ldots,x_{n}\in\mathcal{X}$. The assessment task is to estimate the diversity of generative model $\mathcal{G}$ by measuring the variety of the observed generated data, $x_{1},\ldots x_{n}$. In the following subsections, we will discuss kernel functions and their application to define the VENDI diversity score.
@@ 3.1 Kernel Functions and Matrices
Both the mentioned kernel examples belong to normalized kernels which require $k(x,x)=1$ for every $x$, i.e. the feature map $\phi(x)$ has unit Euclidean norm for every $x$. Given a normalized kernel function, the non-negative eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ points $x_{1},\ldots x_{n}$ will sum up to $1$, which means that they form a probability model.

### v3

# Do Vendi Scores Converge with Finite Samples? Truncated Vendi Score for Finite-Sample Convergence Guarantees

@@ Abstract
Evaluating the diversity of generative models without reference data poses methodological challenges. The reference-free Vendi [Friedman and Dieng, 2023] and RKE [Jalali et al., 2023] scores address this by quantifying the diversity of generated data using matrix-based entropy measures. Among these two, the Vendi score is typically computed via the eigendecomposition of an $n\times n$ kernel matrix constructed from n generated samples. However, the prohibitive computational cost of eigendecomposition for large $n$ often limits the number of samples used to fewer than 20,000. In this paper, we investigate the statistical convergence of the Vendi and RKE scores under restricted sample sizes. We numerically demonstrate that, in general, the Vendi score computed with standard sample sizes below 20,000 may not converge to its asymptotic value under infinite sampling. To address this, we introduce the $t$-truncated Vendi score by truncating the eigenspectrum of the kernel matrix, which is provably guaranteed to converge to its population limit with $n=\mathcal{O}(t)$ samples. We further show that existing Nyström and FKEA approximation methods converge to the asymptotic limit of the truncated Vendi score. In contrast to the Vendi score, we prove that the RKE score enjoys universal convergence guarantees across all kernel functions. We conduct several numerical experiments to illustrate the concentration of Nyström and FKEA computed Vendi scores around the truncated Vendi score, and we analyze how the truncated Vendi and RKE scores correlate with the diversity of image and text data. The code is available at https://github.com/aziksh-ospanov/truncated-vendi.
@@ 1 Introduction
Figure 1: Statistical convergence of Vendi and RKE scores for different sample sizes on ImageNet data: (Left plots) finite-dimension cosine similarity kernel (Right plots) infinite dimension Gaussian kernel with bandwidth $\sigma=30$. The RKE and truncated Vendi scores converged with below 20000 samples, but the Vendi score with Gaussian kernel did not converge.
The increasing use of generative artificial intelligence has underscored the need for accurate evaluation of generative models. In practice, users often have access to multiple generative models trained with different training datasets and algorithms, requiring evaluation methods to identify the most suitable model. The feasibility of a model evaluation approach depends on factors such as the required generated sample size, computational cost, and the availability of reference data. Recent studies on evaluating generative models have introduced assessment methods that relax the requirements on data and computational resources.
Specifically, to enable the evaluation of generative models without reference data, the recent literature has focused on reference-free evaluation scores that remain applicable in the absence of reference samples. The Vendi score [Friedman and Dieng, 2023] is one such reference-free metric that quantifies the diversity of generated data using the entropy of a kernel similarity matrix formulated for the generated samples. Given the sorted eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{n}$ of the normalized matrix $\frac{1}{n}K$ [^1] In general, we consider the trace-normalized kernel matrix $\frac{1}{\mathrm{Tr}(K)}K$, which given $\forall x:\>k(x,x)=1$, reduces to $\frac{1}{n}K$. for the kernel similarity matrix $K=\bigl{[}k(x_{i},x_{j})\bigr{]}_{1\leq i,j\leq n}$ of $n$ generated samples $x_{1},\ldots,x_{n}$, the definition of (order-1) Vendi score is as:
$\mathrm{Vendi}(x_{1},\ldots,x_{n})\,:=\,\exp\Bigl{(}\,\sum_{i=1}^{n}\lambda_{i}\log\frac{1}{\lambda_{i}}\,\Bigr{)}$ (1)
Following conventional definitions in information theory, the Vendi score corresponds to the exponential of the Von Neumann entropy of normalized kernel matrix $\frac{1}{n}K$. More generally, Jalali et al. [2023] define the Rényi Kernel Entropy (RKE) score by applying order-2 Rényi entropy to this matrix, which reduces to the inverse-squared Frobenius norm of the normalized kernel matrix:
$\mathrm{RKE}(x_{1},\ldots,x_{n})\,:=\frac{1}{\Bigl{\|}\frac{1}{n}K\Bigr{\|}^{2}_{F}}$ (2)
Although the Vendi and RKE scores do not require reference samples, their computational cost increases rapidly with the number of generated samples $n$. Specifically, calculating the Vendi score for the $n\times n$ kernel matrix $K$ generally involves an eigendecomposition of $K$, requiring $O(n^{3})$ computations. Therefore, the computational load of Vendi score becomes substantial for a large sample size $n$, and the Vendi score is typically evaluated for sample sizes limited to 20,000. In other words, the Vendi score, as defined in Equation (2), would be computationally infeasible to compute with standard processors for sample sizes greater than a few tens of thousands.
Following the above discussion, a key question that arises is whether the Vendi score estimated from restricted sample sizes (i.e. $n\leq 20000$) has converged to its asymptotic value with infinite samples, which we call the population Vendi. However, the statistical convergence of the Vendi score has not been thoroughly investigated in the literature. In this work, we study the statistical convergence of the Vendi and RKE diversity scores and aim to analyze the concentration of the estimated scores from a limited number of generated samples $n\lessapprox 20000$.
1.1 Our Results on Vendi’s Convergence
@@
We discuss the answer to the Vendi convergence question for two types of kernel functions: 1) kernel functions with a finite feature dimension, e.g. the cosine similarity and polynomial kernels, 2) kernel functions with an infinite feature map such as Gaussian (RBF) kernels. For kernel functions with a finite feature dimension $d$, we theoretically and numerically show that a sample size $n=O(d)$ is sufficient to guarantee convergence to the population Vendi (asymptotic value when $n\rightarrow\infty$). For example, the left plot in Figure 1 shows that in the case of the cosine similarity kernel, the Vendi score on $n$ randomly selected ImageNet [Deng et al., 2009] samples has almost converged as the sample size reaches 5000, where the dimension $d$ (using standard DINOv2 embedding [Oquab et al., 2023]) is 768.
@@
In contrast, our numerical results for kernel functions with an infinite feature map demonstrate that for standard datasets, a sample size bounded by 20,000 could be insufficient for convergence of the Vendi score. For example, the right plot of Figure 1 shows the evolution of the Vendi score with the Gaussian kernel on ImageNet data, and the score continues to grow at a significant rate with 20,000 samples [^2] The heavy computational cost prohibits an empirical evaluation of the sample size required for Vendi’s convergence..
Observing the difference between Vendi score convergence for finite and infinite-dimension kernel functions, a natural question is how to extend the definition of Vendi score from finite to infinite dimension case such that the diversity score would statistically converge in both scenarios. We attempt to address the question by introducing an alternative Vendi statistic, which we call the $t$-truncated Vendi score. The $t$-truncated Vendi score is defined using only the top-$t$ $\lambda_{1}\geq\cdots\geq\lambda_{t}$ eigenvalues of the kernel matrix, where $t$ is an integer hyperparameter. This modified score is defined as
$\mathrm{Truncated}\text{-}\mathrm{Vendi}^{(t)}(x_{1},\ldots,x_{n})=\exp\Bigl{(}\sum_{i=1}^{t}{\lambda^{\scriptscriptstyle\text{trunc}}_{i}}\log\frac{1}{\lambda^{\scriptscriptstyle\text{trunc}}_{i}}\Bigr{)}$
where we shift each of the top-$t$ eigenvalue $\lambda^{\scriptscriptstyle\text{trunc}}_{i}=\lambda_{i}+c$ by the same constant $c=\bigl{(}1-\sum_{i=1}^{t}\lambda_{i}\bigr{)}/t$ to ensure they add up to $1$ and provide a valid probability model. Observe that for a finite kernel dimension $d$ satisfying $d\leq t$, the truncated and original Vendi scores take the same value, because the truncation will have no impact on the eigenvalues. On the other hand, under an infinite kernel dimension, the two scores may take different values.
As a main theoretical result, we prove that a sample size $n=O(t)$ is always enough to estimate the $t$-truncated population Vendi from $n$ empirical samples, regardless of the finiteness of the kernel feature dimension. This result shows that the $t$-truncated Vendi score provides a statistically converging extension of the Vendi score from the finite kernel dimension to the infinite dimension case. To connect the defined $t$-truncated Vendi score to existing computation methods for the original Vendi score, we show that the existing computationally-efficient methods for computing the Vendi score can be viewed as approximations of our defined $t$-truncated Vendi. Specifically, we show that the Nyström method in [Friedman and Dieng, 2023] and the FKEA method proposed by Ospanov et al. [2024b] provide an estimate of the $t$-truncated Vendi.
Figure 2: Computation of the proposed $t$-truncated Vendi score. The kernel similarity matrix eigenspectrum is truncated, and the mass of the truncated tail (excluding the top-$t$ eigenvalues) is uniformly redistributed among the top-$t$ eigenvalues.
1.2 Our Results on RKE’s Convergence
For the RKE score, we prove a universal convergence guarantee that holds for every kernel function. The theoretical guarantee shows that the RKE score, and more generally every order-$\alpha$ entropy score with $\alpha\geq 2$, will converge to its population value within $O(\frac{1}{\sqrt{n}})$ error for $n$ samples. Our theoretical guarantee also transfers to the truncated version of the RKE score. However, note that the truncation of the eigenspectrum becomes unnecessary in the RKE case, since the score enjoys universal convergence guarantees. Figure 1 shows that using both the cosine-similarity and Gaussian kernel functions, the RKE score nearly converges to its limit value with less than 10000 samples.
@@ 1.2 Our Results on RKE’s Convergence
Finally, we present the findings of several numerical experiments to validate our theoretical results on the convergence of Vendi, truncated Vendi, and RKE scores. Our numerical results on standard image, text, and video datasets and generative models indicate that in the case of a finite-dimension kernel map, the Vendi score can converge to its asymptotic limit, in which case, as we explained earlier, the Vendi score is identical to the truncated Vendi. On the other hand, in the case of infinite-dimension Gaussian kernel functions, we numerically observe the growth of the score beyond $n=$10,000. Our numerical results further confirm that the scores computed by Nyström method in [Friedman and Dieng, 2023] and the FKEA method [Ospanov et al., 2024b] provide tight estimations of the population truncated Vendi. The following summarizes this work’s contributions:
  • Analyzing the statistical convergence of Vendi and RKE diversity scores under restricted sample sizes $n\lessapprox 2\times 10^{4}$,
  • Providing numerical evidence on the Vendi score’s lack of convergence for infinite-dimensional kernel functions, e.g. the Gaussian (RBF) kernel,
  • Introducing the truncated Vendi score as a statistically converging extension of the Vendi score from finite to infinite dimension kernel functions,
  • Demonstrating the universal convergence of the RKE diversity score across all kernel functions.
@@ 2 Related Works
Diversity evaluation for generative models Diversity evaluation in generative models can be categorized into two primary types: reference-based and reference-free methods. Reference-based approaches rely on a predefined dataset to assess the diversity of generated data. Metrics such as FID [Heusel et al., 2018], KID and distributed KID [Bińkowski et al., 2018, Wang et al., 2023] measure the distance between the generated data and the reference, while Recall [Sajjadi et al., 2018, Kynkäänniemi et al., 2019] and Coverage [Naeem et al., 2020] evaluate the extent to which the generative model captures existing modes in the reference dataset. Pillutla et al. [2021, 2023] propose MAUVE metric that uses information divergences in a quantized embedding space to measure the gap between generated data and reference distribution. In contrast, the reference-free metrics, Vendi [Friedman and Dieng, 2023] and RKE [Jalali et al., 2023], assign diversity scores based on the eigenvalues of a kernel similarity matrix of the generated data. Jalali et al. [2023] interpret the approach as identifying modes and their frequencies within the generated data followed by entropy calculation for the frequency parameters. The Vendi and RKE scores have been further extended to quantify the diversity of conditional prompt-based generative AI models [Ospanov et al., 2024a, Jalali et al., 2024] and to select generative models in online settings [Rezaei et al., 2024, Hu et al., 2024, 2025]. Also, [Zhang et al., 2024, 2025, Jalali et al., 2025, Gong et al., 2025, Wu et al., 2025] extend the entropic kernel-based scores to measure novelty and embedding dissimilarity. In our work, we specifically focus on the statistical convergence of the vanilla Vendi and RKE scores.
Statistical convergence analysis of kernel matrices’ eigenvalues. The convergence analysis of the eigenvalues of kernel matrices has been studied by several related works. Shawe-Taylor et al. [2005] provide a concentration bound for the eigenvalues of a kernel matrix. We note that the bounds in [Shawe-Taylor et al., 2005] use the expectation of eigenvalues $\mathbb{E}_{m}[\hat{\boldsymbol{\lambda}}(S)]$ for a random dataset $S=(\mathbf{x}_{1},\dots,\mathbf{x}_{m})$ of fixed size $m$ as the center vector in the concentration analysis. However, since eigenvalues are non-linear functions of a matrix, this concentration center vector $\mathbb{E}_{m}[\hat{\boldsymbol{\lambda}}(S)]$ does not match the eigenvalues of the asymptotic kernel matrix as the sample size approaches to infinity. On the other hand, our convergence analysis focuses on the asymptotic eigenvalues with an infinite sample size, which determines the limit value of Vendi scores. In another related work, Bach [2022] discusses a convergence result for the Von-Neumann entropy of kernel matrix. While this result proves a non-asymptotic guarantee on the convergence of the entropy function, the bound may not guarantee convergence at standard sample sizes for computing Vendi scores (less than $10000$ in practice). In our work, we aim to provide convergence guarantees for the finite-dimension and generally truncated Vendi scores with restricted sample sizes.
@@
Efficient computation of matrix-based entropy. Several strategies have been proposed in the literature to reduce the computational complexity of matrix-based entropy calculations, which involve the computation of matrix eigenvalues—a process that scales cubically with the size of the dataset. Dong et al. [2023] propose an efficient algorithm for approximating matrix-based Renyi’s entropy of arbitrary order $\alpha$, which achieves a reduction in computational complexity down to $O(n^{2}sm)$ with $s,m\ll n$. Additionally, kernel matrices can be approximated using low-rank techniques such as incomplete Cholesky decomposition [Fine and Scheinberg, 2001, Bach and Jordan, 2002] or CUR matrix decompositions [Mahoney and Drineas, 2009], which provide substantial computational savings. Pasarkar and Dieng [2024] suggest to leverage Nyström method [Williams and Seeger, 2000] with $m$ components, which results in $O(nm^{2})$ computational complexity. Further reduction in complexity is possible using Random Fourier Features, as suggested by Ospanov et al. [2024b], which allows the computation to scale linearly with $O(n)$ as a function of the dataset size. This work focuses on the latter two methods and the population quantities estimated by them.
@@
Impact of embedding spaces on diversity evaluation. In our image-related experiments, we used the DinoV2 embedding [Oquab et al., 2023], as Stein et al. [2023] demonstrate the alignment of this embedding with human evaluations. We note that the kernel function in the Vendi score can be similarly applied to other embeddings, including the standard InceptionV3[Szegedy et al., 2016] and CLIP embeddings [Radford et al., 2021] as suggested by Kynkäänniemi et al. [2022].
@@ 3 Preliminaries
Consider a generative model $\mathcal{G}$ that generates samples from a probability distribution $P_{X}$. To conduct a reference-free evaluation of the model, we suppose the evaluator has access to $n$ independently generated samples from $P_{X}$, denoted by $x_{1},\ldots,x_{n}\in\mathcal{X}$. The assessment task is to estimate the diversity of generative model $\mathcal{G}$ by measuring the variety of the observed generated data, $x_{1},\ldots x_{n}$. In the following subsections, we will discuss kernel functions and their application to define the Vendi and RKE diversity scores.
@@ 3.1 Kernel Functions and Matrices
Following the standard definition, $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ is called a kernel function if for every integer $n\in\mathbb{N}$ and inputs $x_{1},\ldots,x_{n}\in\mathcal{X}$, the following kernel similarity matrix $K\in\mathbb{R}^{n\times n}$ is positive semi-definite (PSD):
$K=\begin{bmatrix}k(x_{1},x_{1})&\cdots&k(x_{1},x_{n})\\ \vdots&\ddots&\vdots\\ k(x_{n},x_{1})&\cdots&k(x_{n},x_{n})\end{bmatrix}$ (3)
@@
Both the mentioned kernel examples belong to normalized kernels which require $k(x,x)=1$ for every $x$, i.e., the feature map $\phi(x)$ has unit Euclidean norm for every $x$. Given a normalized kernel function, the non-negative eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ points $x_{1},\ldots x_{n}$ will sum up to $1$, i.e., they form a probability model.